a rectangular walkway is 18 feet longer than it is wide. If the area of the walkway is 88 square feet, find the length and the width. (Points :5)

Answer by mananth(16946)

let the width be x
length = x+18
area = L*W
x(x+18)=88
x^2+18x-88=0
x^2+22x-4x-88=0
x(x+22)-4(x+22)=0
(x+22)(x-4)=0
x=-22 OR x=4
width is 4 feet
length = 18+4=22 feet
Ananth